# Dynamic instabilities and turbulence of merged rotating Bose-Einstein condensates

### v1

@@ Abstract
We present the simulation results of merging harmonically confined rotating Bose-Einstein condensates in two dimensions. Merging of the condensate is triggered by positioning the rotation axis at the trap minima and moving both condensates towards each other while slowly ramping their rotation frequency. We analyze the dynamics of the merged condensate by letting them evolve under a single harmonic trap. We systematically investigate the formation of solitonic and vortex structures in the final, unified condensate, considering both non-rotating and rotating initial states. In both cases, merging leads to the formation of solitons that decay into vortex pairs through snake instability and subsequently, these pairs annihilate. Soliton formation and decay-induced phase excitations generate sound waves, more pronounced when the merging time is short. We witness no sound wave generation at sufficiently longer merging times that finally leads to the condensate reaching its ground state. With rotation, we notice off-axis merging (where the rotation axes are not aligned), leading to the distortion and weakening of soliton formation. The incompressible kinetic energy spectrum exhibits a Kolmogorov-like cascade [$E(k)\sim k^{-5/3}$] in the initial stage for merging condensates rotating above a critical frequency and a Vinen-like cascade [$E(k)\sim k^{-1}$] at a later time for all cases. Our findings hold potential significance for atomic interferometry, continuous atomic lasers, and cosmic events related to the remnants of binary neutron-star mergers.
@@ I Introduction
Mergers of Bose-Einstein condensates (BECs) have been the topic of great interest among the ultracold community owing to their wide applications in building up many futuristic tools such as matter-wave interferometry, continuous atom lasers, atmotronics, etc. [1, 2, 3]. Experimentally condensate mergers have been realized in the laboratory and further investigated extensively by Scherer and group [4], who demonstrated a systematic formation of the vortices as a result of the controlled merging between the condensates. This observation played a pivotal role in establishing a one-to-one relation between the dynamical instability of merged condensate along with the appearance of topological defects formed via phase transitions, popularly termed the Kibble-Zurek mechanism [5, 6]. To this end, another successful experiment on the condensate merger was carried out by Chikkatur [7] using optical tweezers, who observed the evaporative loss of atoms in the final state of the condensate and further systematically demonstrated that the randomized initial phases could replenish BECs which further assist in developing the vortex structures in the merged state. Condensate mergers are not restricted to laboratory experiments only; it has been the essential tool in explaining the structure and dynamics of remnants of binary neutron-star merger event GW170817 [8, 9]. For instance, there are some works [10, 11] that suggest that the merger of binary stars that are mainly made of incompressible gluon- quark superfluid cores include several time scales, such as low-energy slow fusion of the core and the fast merger of surrounding envelopes of compressible and dissipative normal matter, which are responsible for the observation of pulsars.
@@
Several works highlight the formation of interference patterns and their properties resulting from the mergers [2, 12, 1, 4, 13, 14, 15]. The study of the merging process and the subsequent formation of interference patterns and solitary waves has been extended to the realms of self-gravitating BECs where fundamental solitons not only form in these head-on collisions but also survive and retain their original shapes [16, 17]. Further studies on collisions of self-gravitating BECs include Paredes’ observation of destructive interference patterns in solitonic collisions [18] and Schwabe’s studies on gravitational cooling and density oscillations in merging of solitonic cores [19]. Recently, Nikolaieva et al. [20] investigated the stability of colliding vortex solitons in self-gravitating BECs for different angular momentum orientations. Similar to previous literature, they also reported the survival of solitons after head-on collisions. Self-gravitating BECs are shown to be excellent candidates for modelling rotating neutron-star interiors and their merging events [21, 9].
@@
The merger between the condensates and the follow-up dynamical state of the merged condensate depends very much on the phases between the condensates before merging. For instance, Stickney and their colleagues [22] demonstrated that the mergers of condensates with opposite phases exhibit excited modes in the merged condensate, leading to the generation of unstable modes. The follow-up works by Mebrahtu [23] and Yi [24] suggested the phase-sensitive excitations and the influence of initial phase differences. Xiong et al. [12] analyzed the effect of the relative initial phase difference on the merging dynamics. Carretero-Gonzalez [25] investigated the role of recombination times on the overall structure and dynamics of the merged condensate. Buchmann et al. [26] numerically extended the experimental study of [4] and also explored the effects of recombination/ramping time on vortex formation in a merged condensate for a given initial relative phase and conclusively demonstrated a possibility of the formation of vortex cores during mergers that very much depends on the choice of phase difference between the condensates. Scott et al. [27] reported the formation of persistent dark solitary waves from interference patterns, which in turn were formed from counterpropagating BEC clouds. These works, however, exclude the role of phase excitations and phase differences in the soliton dynamics. Jo et al. [3] reported the appearance of condensate heating and its influence on atomic numbers using external potential gradients. Apart from this, the dynamics of BECs confined in double-well potentials that exhibit a variety of complex phases have been systematically analyzed [28, 29, 30]. In recent years, the self-bound state, mainly the quantum droplet, has been considered as a prototype model to investigate the different sorts of instabilities arising due to the elastic and inelastic collision between the solitons [31, 32].
@@
Subsequently, in the recent past, there have been several experimental and numerical investigations that focused on the role of dynamical instabilities manifest in the system as a result of merging, which eventually takes the system to a turbulent state. For instance, Kanai et al. [33, 34, 35] considered co-axial mergers of rotating BECs with non-uniform initial phases, where they observed the formation of spiral dark-solitons in both 2D and 3D cases, which facilitated the transfer of angular momentum between condensates. This angular momentum transfer is proportional to the initial angular momentum density difference. Furthermore, they also confirmed the formation of vortices by the decay of solitons via snake instability.
@@
During the past few decades, research on turbulence in 2D condensates has been extensive, encompassing both decaying and forced turbulence [36, 37, 38, 39]. Recently, rotational quantum turbulence has attracted significant interest due to its fascinating properties, contrasting with its non-rotating counterpart [40, 41]. While previous studies have identified dynamical instability in merged Bose-Einstein condensates (BECs), suggesting a potential connection to turbulence, a detailed characterization of turbulence generated through mergers is still lacking. Therefore, further investigation is necessary to understand how the recombination/merging time scales influence the generation of phase excitations and their manifestation as compressible sound waves. Additionally, prior analyses of rotational mergers typically focus on condensates merging and rotating around the same axis. Although Nikolaieva [20] explored mergers in self-gravitating rotating BECs with different angular momentum orientations, the effect of the rotation has not been considered.
@@
In this work, we consider spatially separated, harmonically confined condensates initialized with random phases and rotating at the same frequency. We move them towards each other in real time by updating the respective trap minima and the position of the rotational axis of each condensate along the $x$-axis. Once the condensates merge at the center, we let the merged condensate evolve and analyze its behaviour. We also double the final rotation frequency compared to the initial ones. We investigate the nature of quantum turbulence and its associated incompressible kinetic energy spectra for rotating and non-rotating mergers under different merging times. We highlight the formation of dark solitons and how their dynamics are affected by rotation. Also, we examine the possibility of an adiabatic merger, where phase excitations and the suppression of subsequent formation of compressible sound waves in both rotating and non-rotating cases.
@@ II Mean-field Model and simulation details
We consider a quasi-two-dimensional condensate with strong confinement along the $z$-axis with rotating frequency $\Omega$. The dynamical model describing a system of two condensates is given by the following Gross-Pitaevskii (GP) equation
@@
$\mathrm{i}\frac{\partial\psi}{\partial t}=$ $\left[-\frac{1}{2}\nabla^{2}+V(\mathbf{r},t)+g_{2D}\lvert\psi\rvert^{2}-\Omega(t)L_{z}\right]\psi,$ (1)
@@
where $\psi\equiv\psi(\mathbf{r},t)$ denotes the condensate wave function, with $\mathbf{r}\equiv(x,y)$, $\nabla^{2}$ is the two-dimensional Laplace operator defined as $\nabla^{2}\equiv\partial_{x}^{2}+\partial_{y}^{2}$, and $V(\mathbf{r})$ is the external potential, which includes the harmonic trap along with the central circular barrier. The nonlinear term $g_{2D}=4\pi aN/(\sqrt{2\pi}d_{z})$ represents the interaction strength between the atoms, where $N$ is the total number of atoms, $a$ is the s-wave scattering length, and $d_{z}$ corresponds to the axial width of the trap. Here, $\Omega(t)$ represents the rotational frequency, and $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$ is the $z$ component of angular momentum.
@@
In Eq. (1), the unit length is measured in terms of the transverse harmonic oscillator length $l_{\mbox{ho}}=\sqrt{\hbar/(m\omega)}$, time is measured in units of $\omega^{-1}$, and trap frequency is measured in terms of $\omega$, where, $\omega$ is the transverse trap frequency. For both rotating and non-rotating cases, we fix $N=10^{4}$ and $g_{2D}=400$. For the above choices of parameters, we obtain a scattering length of $a\approx 3.8a_{0}$. The BECs should have a strong axial confinement along $z$ axis so as to restrict the dynamics within the $x-y$ plane by implementing trap frequencies $\omega_{x}\sim 2\pi\times 33\mathrm{Hz}$, $\omega_{y}\sim 2\pi\times 33\mathrm{Hz}$, $\omega_{z}\sim 2\pi\times 1.5\mathrm{kHz}$. This configuration provides a trap frequency of $\omega=\left(\omega_{x}\omega_{y}\omega_{z}\right)^{1/3}\sim 2\pi\times 117.8\mathrm{Hz}$, harmonic oscillator length $l_{\mbox{ho}}\approx 1\mu$m and unit time $\omega^{-1}\approx 1.3$ms. The oscillator length along the $z$ direction would be $l_{z}\approx 0.28\mu$m. Experimentally, for a condensate of ${}^{87}\mathrm{Rb}$ atoms, the desired scattering length can be accessible by tuning the magnetic fields utilizing the Feshbach resonance [42, 43].
@@
To demonstrate the merger of the separated harmonic traps, we utilize the potential model and merging profile prescribed by Buchmann et al. [26]. The potential is given by
$V(\mathbf{r},t)=\frac{1}{2}\left[\gamma(\lvert x\rvert-s(t)l)^{2}+\nu y^{2}\right],$ (2)
where, $\gamma=1$ and $\nu=1$ are trap aspect ratios, $l=6.4$ is the initial position of the trap minimum. The function $s(t)$ determines the merging time $t_{m}$ which is defined as
@@
Initially, at $t=0$, the traps are well-separated with their minima located at ±$\pm l$, where $l$ is chosen large enough to prevent the condensate wavefunctions from overlapping. During the merging process, for $0<t<t_{m}$, we allow the traps to move towards each other. The time evolution of the trap movement is characterized by the positions of the trap minima, denoted by ±$\pm s(t)l$. As the condensates approach each other, the barrier between them reduces, and subsequently, at the end of merging, there is a complete overlap between the two condensates. In the entire duration after the complete merging, i.e., $t>t_{m}$, the condensate remains trapped within a single harmonic well.
To characterize the turbulence generated in the condensate due to the complex vortex dynamics, we perform a detailed analysis of the kinetic energy spectra and corresponding fluxes at different length and time scales. For this analysis, we decompose the kinetic energy into compressible and incompressible components [44] by transforming incompressible and compressible velocity fields into $k$ space integrals via Parseval’s theorem. We use the numerical implementation developed by Bradley and colleagues [45], which evaluates these integrals using the angle-averaged Wiener-Khinchin theorem as given below for the incompressible kinetic energy spectrum

### v3

@@ Abstract
We present the simulation results of merging harmonically confined rotating Bose-Einstein condensates in two dimensions. Merging of the condensate is triggered by positioning the rotation axis at the trap minima and moving both condensates towards each other while slowly ramping their rotation frequency. We analyze the dynamics of the merged condensate by letting them evolve under a single harmonic trap. We systematically investigate the formation of solitonic and vortex structures in the final, unified condensate, considering both non-rotating and rotating initial states. In both cases, merging leads to the formation of solitons that decay into vortex pairs through snake instability, and subsequently, these pairs annihilate. Soliton formation and decay-induced phase excitations generate sound waves, more pronounced when the merging time is short. We witness no sound wave generation at sufficiently longer merging times that finally leads to the condensate reaching its ground state. With rotation, we notice off-axis merging (where the rotation axes are not aligned), leading to the distortion and weakening of soliton formation. The incompressible kinetic energy spectrum exhibits a Kolmogorov-like cascade [$E(k)\sim k^{-5/3}$] in the initial stage for merging condensates rotating above a critical frequency and a Vinen-like cascade [$E(k)\sim k^{-1}$] at a later time for all cases. Our findings hold potential significance for atomic interferometry, continuous atomic lasers, and quantum sensing applications.
@@ I Introduction
Mergers of Bose-Einstein condensates (BECs) have been the topic of great interest among the ultracold community owing to their wide applications in building up many futuristic tools such as matter-wave interferometry, continuous atom lasers, atmotronics, etc.Bloch, Hänsch, and Esslinger (2006); Andrews et al. (1997); Jo et al. (2007). Experimentally condensate mergers have been realized in the laboratory and further investigated extensively by Scherer and groupScherer et al. (2007), who demonstrated a systematic formation of the vortices as a result of the controlled merging between the condensates. This observation played a pivotal role in establishing a one-to-one relation between the dynamical instability of merged condensate along with the appearance of topological defects formed via phase transitions, popularly termed the Kibble-Zurek mechanismKibble (1976); Zurek (1985). To this end, another successful experiment on the condensate merger was carried out by ChikkaturChikkatur et al. (2002) using optical tweezers, who observed the evaporative loss of atoms in the final state of the condensate and further systematically demonstrated that the randomized initial phases could replenish BECs which further assist in developing the vortex structures in the merged state.
@@
Several works highlight the formation of interference patterns and their properties resulting from the mergersAndrews et al. (1997); Xiong, Yang, and Benedict (2013); Bloch, Hänsch, and Esslinger (2006); Scherer et al. (2007); Shin et al. (2004); Stock et al. (2005); Li et al. (2024). The study of the merging process and the subsequent formation of interference patterns and solitary waves has been extended to the realms of self-gravitating BECs where fundamental solitons not only form in these head-on collisions but also survive and retain their original shapesChoi (2002); Cotner (2016). Further studies on collisions of self-gravitating BECs include Paredes’ observation of destructive interference patterns in solitonic collisionsParedes and Michinel (2016) and Schwabe’s studies on gravitational cooling and density oscillations in merging of solitonic coresSchwabe, Niemeyer, and Engels (2016). Recently, Nikolaieva et al.Nikolaieva et al. (2023) investigated the stability of colliding vortex solitons in self-gravitating BECs for different orientations of angular momentum. Similar to previous literature, they also reported the survival of solitons after head-on collisions.
@@
The merger between the condensates and the follow-up dynamical state of the merged condensate depends very much on the phases between the condensates before merging. For instance, Stickney and colleaguesStickney and Zozulya (2003) demonstrated that the mergers of condensates with opposite phases exhibit excited modes in the merged condensate, leading to the generation of unstable modes. The follow-up works by MebrahtuMebrahtu, Sanpera, and Lewenstein (2006) and YiYi and Duan (2005) suggested the phase-sensitive excitations and the influence of initial phase differences. Xiong et al.Xiong, Yang, and Benedict (2013) analyzed the effect of the relative initial phase difference on the merging dynamics. Carretero-GonzalezCarretero-González et al. (2008) investigated the role of recombination times on the overall structure and dynamics of the merged condensate. Buchmann et al.Buchmann, Nikolopoulos, and Lambropoulos (2009) numerically extended the experimental study of Scherer et al.Scherer et al. (2007) and also explored the effects of recombination/ramping time on vortex formation in a merged condensate for a given initial relative phase and conclusively demonstrated a possibility of the formation of vortex cores during mergers that very much depends on the choice of phase difference between the condensates. Scott et al.Scott et al. (2005) reported the formation of persistent dark solitary waves from interference patterns, which in turn were formed from counterpropagating BEC clouds. These works, however, exclude the role of phase excitations and phase differences in the soliton dynamics. Jo et al.Jo et al. (2007) reported the appearance of condensate heating and its influence on atomic numbers using external potential gradients. Apart from this, the dynamics of BECs confined in double-well potentials that exhibit a variety of complex phases have been systematically analyzedAdhikari (2014); Roy, Chakrabarti, and Trombettoni (2022); Spagnolli et al. (2017). In recent years, the self-bound state, mainly the quantum droplet, has been considered as a prototype model to investigate the different sorts of instabilities arising due to the elastic and inelastic collision between the solitonsFerioli et al. (2019); Gangwar et al. (2022).
@@
Subsequently, in the recent past, there have been several experimental and numerical investigations that focused on the role of dynamical instabilities manifest in the system as a result of merging, which eventually takes the system to a turbulent state. For instance, Kanai et al.Kanai, Guo, and Tsubota (2019); Kanai et al. (2020); Kanai, Guo, and Tsubota (2018) considered co-axial mergers of rotating BECs with non-uniform initial phases, where they observed the formation of spiral dark-solitons in both 2D and 3D cases, which facilitated the transfer of angular momentum between condensates. This angular momentum transfer is proportional to the initial angular momentum density difference. Furthermore, they also confirmed the formation of vortices by the decay of solitons via snake instability. A closely related work on turbulence aspects is the recent observation of strong quantum turbulence (SQT) in BECs by Middleton-Spencer et al.Middleton-Spencer et al. (2023). They characterized SQT using the fragmented density fluctuations and non-homogeneously distributed vortex loops. SQT was achieved by oscillating the external potential, nucleating solitons, and generating density waves. Notably, Escartin et al. previously observed turbulence in superfluid helium dropletsEscartín et al. (2019), reporting Kolmogorov scaling in the incompressible kinetic spectra due to vortex-antivortex ring nucleation and subsequent weak surface turbulence.
@@
During the past few decades, research on turbulence in 2D condensates has been extensive, encompassing both decaying and forced turbulenceNumasato and Tsubota (2009a, b); Numasato, Tsubota, and L’vov (2010); Das, Mukherjee, and Majumder (2022). Recently, rotational quantum turbulence has attracted significant interest due to its fascinating properties, contrasting with its non-rotating counterpartEstrada, Brachet, and Mininni (2022); Müller et al. (2020). While previous studies have identified dynamical instability in merged Bose-Einstein condensates (BECs), suggesting a potential connection to turbulence, a detailed characterization of turbulence generated through mergers is still lacking. Therefore, further investigation is necessary to understand how the recombination/merging time scales influence the generation of phase excitations and their manifestation as compressible sound waves.
@@
In this work, we consider spatially separated, harmonically confined condensates initialized with random phases and rotating at the same frequency. We move them towards each other in real time by updating the respective trap minima and the position of the rotational axis of each condensate along the $x$-axis. Once the condensates merge at the center, we let the merged condensate evolve and analyze its behaviour. We also double the final rotation frequency compared to the initial ones. We investigate the nature of quantum turbulence and its associated incompressible kinetic energy spectra for rotating and non-rotating mergers under different merging times. We highlight the formation of dark solitons and how their dynamics are affected by rotation. Also, we examine the possibility of an adiabatic merger, where phase excitations and the formation of compressible sound waves are suppressed in both rotating and non-rotating cases.
@@ II Mean-field Model and simulation details
We consider a quasi-two-dimensional condensate with strong confinement along the $z$-axis with a time-dependent rotating frequency $\Omega_{t}(t)$. The dynamical model describing a system of two condensates is given by the following Gross-Pitaevskii (GP) equation in the rotating frame of reference
@@
$\mathrm{i}\frac{\partial\psi}{\partial t}=$ $\left[-\frac{1}{2}\nabla^{2}+V(\mathbf{r},t)+g_{2D}\lvert\psi\rvert^{2}-\Omega_{t}(t)\hat{L}_{z}(t)\right]\psi,$ (1)
@@
where $\psi\equiv\psi(\mathbf{r},t)$ denotes the condensate wave function, with $\mathbf{r}\equiv(x,y)$, $\nabla^{2}$ is the two-dimensional Laplace operator defined as $\nabla^{2}\equiv\partial_{x}^{2}+\partial_{y}^{2}$, and $V(\mathbf{r})$ is the external potential, which includes the harmonic trap along with the central circular barrier. The nonlinear term $g_{2D}=4\pi aN/(\sqrt{2\pi}d_{z})$ represents the interaction strength between the atoms, where $N$ is the total number of atoms, $a$ is the s-wave scattering length, and $d_{z}$ corresponds to the axial width of the trap.
@@
In Eq. (1), the unit length is measured in terms of the transverse harmonic oscillator length $l_{\mbox{ho}}=\sqrt{\hbar/(m\omega)}$, time is measured in units of $\omega^{-1}$, where, $\omega$ is the transverse trap frequency. For both rotating and non-rotating cases, we fix $N=10^{4}$ and $g_{2D}=400$. For the above choices of parameters, we obtain a scattering length of $a\approx 3.8a_{0}$. The BECs should have a strong axial confinement along $z$ axis so as to restrict the dynamics within the $x-y$ plane by implementing trap frequencies $\omega_{x}\sim 2\pi\times 33\mathrm{Hz}$, $\omega_{y}\sim 2\pi\times 33\mathrm{Hz}$, $\omega_{z}\sim 2\pi\times 1.5\mathrm{kHz}$. This configuration provides a trap frequency of $\omega=\left(\omega_{x}\omega_{y}\omega_{z}\right)^{1/3}\sim 2\pi\times 117.8\mathrm{Hz}$, harmonic oscillator length $l_{\mbox{ho}}\approx 1\mu$m and unit time $\omega^{-1}\approx 1.3$ms. The oscillator length along the $z$ direction would be $l_{z}\approx 0.28\mu$m. Experimentally, for a condensate of ${}^{87}\mathrm{Rb}$ atoms, the desired scattering length can be accessible by tuning the magnetic fields utilizing the Feshbach resonanceMarcelis et al. (2004); Bauer et al. (2009).
@@
To demonstrate the merger of the separated harmonic traps, we utilize the potential model and merging profile prescribed by Buchmann et al.Buchmann, Nikolopoulos, and Lambropoulos (2009). The potential is given by
$V(\mathbf{r},t)=\frac{1}{2}\left[\gamma^{2}(\lvert x\rvert-s(t)l)^{2}+\nu^{2}y^{2}\right],$ (2)
where, $\gamma=1$ and $\nu=1$ are trap aspect ratios, $l=6.4$ is the initial position of the trap minimum. Apart from moving the traps, it is also important to move the positions of the rotational axes and also ramp the individual frequencies of the condensates as they approach each other which can be done by modifying the angular momentum operator $\hat{L}_{z}(t)$ and the rotating frequency $\Omega_{t}(t)$, respectively.
$\Omega_{t}(t)=\Omega_{0}(2-s(t))$ (3)
$\hat{L}_{z}(t)=\begin{cases}\left[y\partial_{x}-\left(x-s(t)l\right)\partial_{y}\right],&\mathrm{if}\,x\geq 0\\ \left[y\partial_{x}-\left(x+s(t)l\right)\partial_{y}\right],&\mathrm{if}\,x<0\\ \end{cases}$ (4)
From these equations, one can see that the rotation axes are positioned at the trap minima and are moved simultaneously in real time. When the rotating axes of the two separated condensates approach each other, their rotating frequencies accelerate. Upon complete overlap of the trap and axes, the final rotation frequency reaches $\Omega_{f}$. To ensure a seamless merger of the condensates, we have consistently assumed a final rotation frequency twice that of the initial rotation frequency, i.e., $\Omega_{f}=2\Omega_{0}$.
The function $s(t)$ determines the time $t_{m}$ taken to perform the merger and ramp the individual rotation frequencies which are defined as
@@
At $t=0$, the traps are well-separated, each rotating at a frequency $\Omega_{0}$ and their minima are located at ±$\pm l$. During the merging process, for $0<t<t_{m}$, we allow the traps and the positions of their rotation axes to move towards each other. The time evolution of the trap and the axes movement is characterized by the positions of the trap minima, denoted by ±$\pm s(t)l$. The schematic diagrams in Fig. 1 display the contours and cross-sections of the potential during the merging process.
Figure 1: Contour plots of the potential $V(\mathbf{r},t)$ during the merger process: (a) $t=0$, (b) $0<t<t_{m}$ and (c) $t\geq t_{m}$. At $t=0$, the condensates are well separated and rotate at individual frequencies $\Omega_{0}$. As the merger progresses ($0<t<t_{m}$), the potential barrier between them gradually decreases, and their respective angular frequencies become $\Omega_{t}$ as determined by Eq. (3). Once the merger is completed ($t>t_{m}$), the condensate is confined to a single harmonic well rotating at a frequency of $\Omega_{f}=2\Omega_{0}$.
As the condensates approach each other, the barrier between them reduces, and their respective frequencies increase and, subsequently, at the end of merging, there is a complete overlap between the two condensates now rotating with a frequency of $2\Omega_{0}$. In the entire duration after the complete merging, i.e., $t>t_{m}$, the condensate remains trapped within a single harmonic well and continues to rotate at $2\Omega_{0}$. It is also noteworthy that for sufficiently large separations in the rotating case, the condensates spiral towards each other. We have carefully selected the parameters to minimize and avoid irrelevant dynamics in our studies. For example, we limited the initial individual rotating frequencies to $\Omega_{0}\lesssim 0.5$ to ensure that the final rotating frequency ($\Omega_{f}$) remained below unity. Exceeding this threshold could have caused the merged condensate to escape the harmonic trap. In addition, the nonlinearity strength ($g_{2D}$) was chosen to be sufficient for vortex induction and generation of turbulence. We have also ensured identical initial rotation frequencies for both the condensate and a separation ($l$) between them that was large enough to prevent wave function overlap but not so large as to hinder spiralling effects at shorter times.
To characterize the turbulence generated in the condensate due to the complex vortex dynamics, we perform a detailed analysis of the kinetic energy spectra and corresponding fluxes at different length and time scales. For this analysis, we decompose the kinetic energy into compressible and incompressible componentsNore, Abid, and Brachet (1997) by transforming the incompressible and compressible velocity fields into $k$ space integrals via Parseval’s theorem.
Figure 2: Snapshots of the condensate density at (a) $t=0$, and at different instances for merging times $t_{m}=4$ [(b) - (e)] and $t_{m}=6$ [(f)-(j)]. (b) Formation of interference fringes alongside a central soliton, (c)-(d) shows the formation of vortex-antivortex pairs, and (e) Snapshot after the vortex pair has annihilated. (f) Formation of interference fringes, (g) soliton with significant bending, (h)-(i) formation of vortex pairs, and (j) subsequent annihilation. The anti-vortices and vortices are marked with cross ($\times$) and circle ($\circ$) symbols, respectively.
We use the numerical implementation developed by Bradley and colleaguesBradley et al. (2022), which evaluates these integrals using the angle-averaged Wiener-Khinchin theorem, as given below for the incompressible kinetic energy spectrum